OpenStudy (anonymous):

Find the volume of a right pyramid if the base is a regular pentagon such that its side length is 16ft and the perpendicular distance from its center to one of its sides is 11 ft. The height of the pyramid is 19.2ft.

OpenStudy (anonymous):

area of pentagon = 5 * area of triangle = 5 * 11*8 = 440 sq ft volume of the pyramid = 1/3 * height * area of base = 1/3 * 19.2 * 440 = 2816 cu ft

OpenStudy (anonymous):

is it (5/6)*16*11*19.2 = 2816 cu. ft

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks Sheg and jimmyrep!